OpenStudy (anonymous):

find the value of x..given tht x^4-15x^2-34=0

OpenStudy (anonymous):

Let a=x^2 then, a^2 -15a -34=0 Now, find the value of a then find x

OpenStudy (anonymous):

15x^2 - 34 = 0?

OpenStudy (anonymous):

u just change x to a

OpenStudy (agent0smith):

No, replace x^2 with a. So it becomes \[a^2 -15a -34=0 \] then you can factorize: (a-17)(a+2)=0 so a = 17 and a = -2. And remember a=x^2.

OpenStudy (agent0smith):

Looks like you should get four solutions, two real and two imaginary.